### v2

@@ 4 Discussion
Kinetics: The kinetic analysis revealed critical insights into the biomechanical demands of walking on different terrains. For kinematic profiles such as joint moments compared with on solid ground, sand walking locomotion required less hip extensor action until the mid-stance to control knee flexion and the forward rotation of the upper body. The longitudinal force shown in Fig. 5 has a small backward component at the hip joint which requires small hip extensors consequently. However, compared to solid ground, walking on sand demanded greater hip flexor to pull the thigh forward and upward at the latter half of the stance phase (including the double stance) and get ready for the knee extension and swing. This indicates large energy consumption to compensate for momentum loss at the stance leg. The distinct patterns in GRFs and joint moments suggest that walking on sand necessitates a reorientation of muscular efforts and joint mechanics. This reorientation is a complex interplay between maintaining stability, managing energy efficiency, and adapting to the terrain’s physical characteristics.
@@
This study represents a significant advanced step in analyzing biomechanics during human locomotion across varying terrain types, specifically solid ground and sand. It is the first study that reports the GRFs and joint moments for humans walking on yielding terrains such as sand. The locomotion dataset provided by this study also generates new knowledge and enables to use wearable assistive devices and IMU-based activity recognition for human locomotion on yielding terrains. The kinematic analysis of the lower limbs aligns with existing studies, such as van den Berg et al. (2017). In contrast to Xu et al. (2015), which posited an experimental design that minimizes force dissipation in the sand, the calibration results detailed in A indicate a discrepancy between forces at the sand surface and the GRFs measured by the embedded force plate. Despite requiring calibration for magnitude, the discussed longitudinal and vertical GRFs in this work are in agreement in pattern to those reported by Xu et al. (2015). Contrary to Svenningsen et al. (2019), which found reduced walking speed and stride length on sand, the subjects in this study exhibited increased pace, longer strides, and greater vertical COM displacement. This discrepancy is attributed to the consistent starting conditions for all trials, eliminating bias from self-selected walking strategies per terrain. The methodology in this work thus captures the actual transition between gaits when moving from solid ground to sand. The adaptive locomotion on the sand can be ascribed to the inherent compensatory mechanisms of the human motor system by adjusting stride and body mechanics to maintain balance and stability on the less stable sand surface.
@@
While acknowledging some study limitations, their impact on our findings is not excessively significant. The study’s design permitted subjects to walk at their preferred pace, leading to variability in walking speeds, diverging from other studies like Camargo et al. (2021) focused on locomotion velocity effects. This variability might influence biomechanical data consistency. Additionally, the research scope was limited to straight walking on level sand and solid ground, excluding complex movements like ascending, descending, and turning on granular terrains. Such limitations might restrict a comprehensive understanding of biomechanical adaptations in gait activities. The dataset lacks other lower limb biomechanics, such as EMG signals, important for analyzing muscle activation patterns. Future research should include diverse locomotive modes and participant demographics, using an array of sensors for gait analysis on different terrains to further enhance our understanding of human locomotion on granular terrains.
@@ 5 Conclusion
This study has contributed to the understanding of human walking locomotion on granular terrain by providing a detailed biomechanical analysis of gait adaptations. The findings were derived from comprehensive biomechanical and wearable sensor data from $20$ able-bodied adults and highlighted the intricate adjustments in stride, joint mechanics, and ground reaction forces necessary for efficient movement on challenging surfaces such as sand. These insights offered valuable implications for the development of advanced assistive devices and responsive robotic systems. We also provided and shared open-source kinematic and kinetic dataset and a comparison of human walking on solid and sand surfaces. Despite certain limitations such as the variability in walking speeds and the focus on able-bodied individuals, this research paved the way for future studies to explore a broader range of locomotion conditions and participant demographics and therefore, enhance the applicability of wearable sensing and assistive technologies in diverse environmental settings.
CRediT Authorship Contribution Statement
Chunchu Zhu: Conceptualization, Formal analysis, Methodology, Investigation, Experiments, Software, Writing – original draft. Xunjie Chen: Conceptualization, Formal analysis, Methodology, Investigation, Experiments, Software, Writing – original draft. Jingang Yi: Project administration, Supervision, Resources, Conceptualization, Writing – revision & editing.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
@@ Acknowledgments
This study was partially sponsored by the US National Science Foundation under award CMMI-2222880. The authors thank Mr. Aditya Anikode of Rutgers University for his help in constructing the experimental setup in this study.
@@ Appendix A Force Calibration
Figure 1: Experimental setup for force calibration.
The force plate and sandbox configurations in the experiment were similar to that in Xu et al. (2015). However, the dissipation of the force under this design cannot be negligible. During the walking experiments, the force plate was embedded beneath a layer of sand with a depth of $14$ cm. The GRFs observed at the sand’s surface inherently differed from those recorded by the force plate because of the deformation characteristics and energy dissipation properties of sand substrates. Consequently, it is necessary to conduct force calibration to obtain accurate GRFs from the force plate.
As shown in Fig. 1, we designed and built a sand box with a force plate positioned beneath the sand layer. The calibration setup consists of a lever mechanism, featuring a force/torque load cell (ATI mini45) at one end and a vertically adjustable displacement lift platform at the other end. By raising the platform and leveraging the mechanical advantage of the bearing structure, the compression plate can firmly generate a controlled vertical force onto the sand’s surface. The magnitude of the applied force was recorded by the load cell and was synchronized with the force plate readings. This setup allowed us to establish a correlation between the force exerted at the surface and the force recorded at the bottom, under varying sand thickness conditions.
@@
Figure. 2 shows the force calibration results of a total of fifteen sand thicknesses. The force measured beneath the sand was indeed smaller than that at the surface. Overall, the force ratio $\zeta$ decreases as the sand layer becomes thicker. A significant ratio drop is observed when the sand thickness rises from $7$ to $8$ cm. In summary, this correction force ratio was used to re-calculate the actual GRFs from the force plate measurements. For the subject walking experiments, we used the ratio $\zeta=0.81$ for the $14$-cm sand depth, and therefore, the real vertical ground reaction force $F_{z}^{s}=F_{z}^{b}/\zeta$, where $F_{z}^{b}$ is the force plate measurements in the vertical direction.
Figure 2: The force calibration ratio results at different sand thickness conditions.
@@ Appendix B Inverse Dynamics
In this section, we discuss the calculation of the joint moments in the sagittal plane based on inverse dynamics. The formulations for ankle, knee, and hip joint moments are derived and articulated using Newtonian-Euler formulation specific to the foot, shank, and thigh segments, respectively. Similar derivations can be found in the support materials from the work of Shamaei et al. (2013).
As shown in Fig. 1, the leg consists of three segments, namely, the thigh, shank, and foot segment. For each segment, there are two joints, proximal and distal joints. To obtain the joint moments of each segment, the Newtonian-Euler method was used. Since the procedure is consistent for each segment, we only present the derivation of the proximal and distal joint moments for the foot segment as a brief example.
@@
For the foot segment, the proximal joint is the ankle joint and the distal one is the toe. Here, we consider that the COP is located near the center of the line connecting the toe and heel marker as a simplification assumption, and the ground reaction force vector $\boldsymbol{F}_{G}$ and moment vector $\boldsymbol{M}_{G}$ are applied at the COP point. First, $\boldsymbol{F}_{G}$ and $\boldsymbol{M}_{G}$ should be transferred to the distal joint (toe) such that
@@
Next, according to the Newtonian method, the forces on the foot segment should satisfy $\Sigma\boldsymbol{F}=m_{f}\boldsymbol{a}_{f}$, where $m_{f}$ is the mass of the foot segment and $\boldsymbol{a}_{f}$ is the acceleration of the foot. Therefore, the force at the proximal joint (i.e., the ankle) is
@@
For the knee and hip joint moments, we follow the same process as we treat the foot segment. For instance, we take the ankle and knee joint moments as the distal and proximal joint moments, respectively, for the shank segment, and the knee and hip joint moments are then treated as the distal and proximal joint moments, respectively, for the thigh segment. Then, the knee and hip moments are calculated respectively as
$\boldsymbol{M}_{P}^{S}=$ $-\boldsymbol{M}_{G}-\left(\boldsymbol{r}+l_{f}\boldsymbol{e}_{f}+l_{s}\boldsymbol{e}_{s}\right)\times\boldsymbol{F}_{G}+I_{f}\boldsymbol{\dot{\omega}}_{f}+I_{s}\boldsymbol{\dot{\omega}}_{s}$ (4)
$-\left(l_{p}^{f}\boldsymbol{e}_{f}+l_{s}\boldsymbol{e}_{s}\right)\times m_{f}\left(\boldsymbol{a}_{f}+g\boldsymbol{e}_{z}\right)-l_{p}^{s}\boldsymbol{e}_{s}\times m_{s}\left(\boldsymbol{a}_{s}+g\boldsymbol{e}_{z}\right),$
@@
$\boldsymbol{M}_{P}^{T}=$ $-\left(\boldsymbol{r}+l_{f}\boldsymbol{e}_{f}+l_{s}\boldsymbol{e}_{s}+l_{t}\boldsymbol{e}_{t}\right)\times\boldsymbol{F}_{G}-\left(l_{p}^{s}\boldsymbol{e}_{s}+l_{t}\boldsymbol{e}_{t}\right)\times m_{s}\left(\boldsymbol{a}_{s}+g\boldsymbol{e}_{z}\right)$ (5)
$-\left(l_{p}^{f}\boldsymbol{e}_{f}+l_{s}\boldsymbol{e}_{s}+l_{t}\boldsymbol{e}_{t}\right)\times m_{f}\left(\boldsymbol{a}_{f}+g\boldsymbol{e}_{z}\right)+I_{f}\boldsymbol{\dot{\omega}}_{f}+I_{s}\boldsymbol{\dot{\omega}}_{s}+I_{t}\boldsymbol{\dot{\omega}}_{t}-\boldsymbol{M}_{G},$
@@
The formulations (3), (4), and (5) are consistent for both the stance leg and swing leg. For the swing leg, the ground reaction forces and moments are zero, i.e., $\boldsymbol{F}_{G}=\boldsymbol{0}$ and $\boldsymbol{M}_{G}=\boldsymbol{0}$. Translation accelerations and angular accelerations of the leg segments can be extracted from the optical marker measurements. The corresponding anthropometry information of the subject such as segment length, mass, and location of COM can be found in Chapter 4 of Winter (2009). We used estimate ratios for the joint moment calculation in this study.

### v3

@@ 4.2 Kinetics Characteristics
The kinetic analysis revealed critical insights into the biomechanical demands of walking on different terrains. For kinematic profiles such as joint moments, compared with on solid ground, sand walking locomotion required less hip extensor action until the mid-stance to control knee flexion and the forward rotation of the upper body. The longitudinal force shown in Fig. 5 has a small backward component at the hip joint, which requires small hip extensors consequently. However, compared to solid ground, walking on sand demanded greater hip flexor to pull the thigh forward and upward at the latter half of the stance phase (including the double stance) and to get ready for the knee extension and swing. This indicates large energy consumption to compensate for momentum loss at the stance leg on sand. The distinct patterns in GRFs and joint moments suggest that walking on sand necessitate a reorientation of muscular efforts and joint mechanics. This reorientation is a complex interplay between maintaining stability, managing energy efficiency, and adapting to the terrain’s physical characteristics.
The differences in GRFs between sand and solid ground highlight the significant impact of terrain type on the forces exerted by the foot during walking. The relatively small normalized magnitude of the maximum longitudinal force on sand suggests that the yielding nature of sand reduces the ability to generate propulsive and braking forces, possibly due to the dissipation of energy as the foot interacts with the deformable surface. In contrast, the vertical force on sandy terrain exhibits a higher peak during the heel strike phase and a different double-hump pattern compared to solid ground, indicating that the deformable surface significantly alters the loading pattern during the stance phase. The increased variability and different magnitudes of GRFs on sand also suggest that walking on this surface requires adapted biomechanical strategies to maintain stability and forward progression. These adaptations could involve changes in muscle activation patterns, joint stiffness, and overall gait mechanics to ensure efficient and safe locomotion on the challenging surface.
The knee biomechanical profile potentially provides a tool for assistive device controllers to adjust strategies accordingly. By leveraging joint torque and angle data, assistive controllers can provide phase-specific assistance that aligns with the natural gait cycle, enhancing movement synergy. For example, using the knee exoskeleton in [23], we can further incorporate the active knee biomechanics profile with respect to knee flexion/extension. Adaptive real-time biomechanical feedback control can dynamically adjust assistance on varying terrains and, therefore, improve stability and energy efficiency for the user. This data-driven approach promises to refine the integration of exoskeletons with human locomotion and potentially enhance assistive device functionality [21].
@@
This study represents an advanced step in analyzing biomechanics during human locomotion across varying terrain types, specifically solid ground and sand. It is the first study that reports the GRFs and joint moments for humans walking on yielding terrains such as sand. The locomotion dataset provided by this study also generates new knowledge and enables to use wearable assistive devices and IMU-based activity recognition for human locomotion on yielding terrains. The kinematic analysis of the lower limbs aligns with existing studies, such as [24]. In contrast to [10], which posited an experimental design that minimizes force dissipation in the sand, the calibration results detailed in Appendix A indicate a discrepancy between forces at the sand surface and the GRFs measured by the embedded force plate. Despite requiring calibration for magnitude, the reported longitudinal and vertical GRFs in this work are in agreement in pattern to those reported by [10]. Contrary to [3], which found reduced walking speed and stride length on sand, the participants in this study exhibited increased pace, longer strides, and greater vertical COM displacement. This discrepancy is attributed to the consistent starting conditions for all trials, eliminating bias from self-selected walking strategies per terrain.
The methodology in this work thus captures the actual transition between gaits when moving from solid ground to sand. The adaptive locomotion on the sand can be ascribed to the inherent compensatory mechanisms of the human motor system by adjusting stride and body mechanics to maintain balance and stability on the less stable sand surface. The 7.5-meter walkway used in this study, although relatively short, provides valuable insights into the immediate adaptations and biomechanical changes that occur when transitioning between solid ground and sand. The platform design allows for the capture of at least one complete nature gait cycle on each terrain type, enabling the analysis of important biomechanical parameters. This study design is particularly useful for understanding the strategies employed by the human locomotion system when confronted with sudden changes in surface properties, providing valuable insights into the mechanisms of gait adaptation on varying surfaces.
@@
While acknowledging some study limitations, their impact on our findings is not significant. The study’s design permitted participants to walk at their preferred pace, leading to variability in walking speeds, diverging from other studies such as [1] focused on locomotion velocity effects. This variability might influence biomechanical data consistency. In [10], the results showed the magnitude of COP changes on level ground in the anterior-posterior (AP) direction is within 4 mm when the sand deformation is 20 mm. Therefore, for simplicity, we take the COP position from the GRF data. Additionally, the research scope was limited to straight walking on level sand and solid ground, excluding complex movements like ascending, descending, and turning on granular terrains. Such limitations might restrict a comprehensive understanding of biomechanical adaptations in gait activities. The dataset lacks other lower limb biomechanics, such as EMG signals, important for analyzing muscle activation patterns. Future research should include diverse locomotive modes and participant demographics, using an array of sensors for gait analysis on different terrains to further enhance our understanding of human locomotion on granular terrains.
@@ 5 Conclusion
This study has contributed to the understanding of human walking locomotion on granular terrain by providing a detailed biomechanical analysis of gait adaptations. The findings were derived from comprehensive biomechanical and wearable sensor data from $20$ able-bodied adults and highlighted the intricate adjustments in stride, joint mechanics, and ground reaction forces necessary for efficient movement on challenging surfaces such as sand. These insights offered valuable implications for the development of advanced assistive devices and responsive robotic systems. We also provided and shared open-source kinematic and kinetic dataset and a comparison of human walking on solid and sand surfaces. This research work provided biomechanics dataset and methodology for future studies to explore a broader range of locomotion conditions and participant demographics and therefore, enhance the applicability of wearable sensing and assistive technologies in diverse environmental settings.
@@ Acknowledgments
The authors thank Mr. Aditya Anikode of Rutgers University for his help in constructing the experimental setup in this study. The work was partially supported by US National Science Foundation (NSF) under award CMMI-2222880.
Appendices
@@ Appendix A Force Calibration
The force plate and sandbox configurations in the experiment were similar to that in [10]. However, the dissipation of the force under this design cannot be negligible. During the walking experiments, the force plate was embedded beneath a layer of sand with a depth of $14$ cm. The GRFs observed at the sand’s surface inherently differed from those recorded by the force plate because of the deformation characteristics and energy dissipation properties of sand substrates. Consequently, it is necessary to conduct force calibration to obtain accurate GRFs from the force plate.
As shown in Fig. 8, we designed and built a sand box with a force plate positioned beneath the sand layer. The calibration setup consists of a lever mechanism, featuring a force/torque load cell (ATI mini45) at one end and a vertically adjustable displacement lift platform at the other end. By raising the platform and leveraging the mechanical advantage of the bearing structure, the compression plate firmly generated a controlled vertical force onto the sand’s surface. The magnitude of the applied force was recorded by the load cell and data collection was synchronized with the force plate readings. This setup allowed us to establish a correlation between the force exerted at the surface and the force recorded by the embedded force plate under varying sand thickness conditions.
Figure 8: Experimental setup for force calibration on sand terrain.
@@
Figure 9 shows the force calibration results. The force measured beneath the sand was indeed smaller than that at the surface. Overall, the force ratio $\zeta$ decreases as the sand layer becomes thicker. A significant ratio drop is observed when the sand thickness rises from $7$ to $8$ cm. In summary, this correction force ratio was used to re-calculate the actual GRFs from the force plate measurements. For the participant walking experiments, we used the ratio $\zeta=0.81$ for the $14$-cm sand depth, and therefore, the real vertical ground reaction force $F_{z}^{s}=F_{z}^{b}/\zeta$, where $F_{z}^{b}$ is the force plate measurements in the vertical direction.
Figure 9: The force calibration ratio results for force plate embedded at different sand depths.
@@ Appendix B Inverse Dynamics
In this section, we present the calculation of the joint moments in the sagittal plane based on inverse dynamics. The formulations for ankle, knee, and hip joint moments are derived and articulated using Newtonian-Euler formulation specific to the foot, shank, and thigh segments, respectively. Similar derivations can be found in the support materials from the work of [22].
As shown in Fig. 10, the leg consists of three segments, namely, the thigh, shank, and foot segment. For each segment, there are two joints, proximal and distal joints. To obtain the joint moments of each segment, the Newtonian-Euler method is used. Since the procedure is consistent for each segment, we only present the derivation of the proximal and distal joint moments for the foot segment as a brief example. For the foot segment, the proximal joint is the ankle joint and the distal one is the toe, and the ground reaction force vector $\boldsymbol{F}_{G}$ and moment vector $\boldsymbol{M}_{G}$ are applied at the COP point.
@@
First, $\boldsymbol{F}_{G}$ and $\boldsymbol{M}_{G}$ should be transferred to the distal joint (toe) such that
@@
Next, the forces on the foot segment should satisfy $\Sigma\boldsymbol{F}=m_{f}\boldsymbol{a}_{f}$, where $m_{f}$ is the mass of the foot segment and $\boldsymbol{a}_{f}$ is the acceleration of the foot. Therefore, the force at the proximal joint (i.e., the ankle) is
@@
For the knee and hip joint moments, we follow the same process as we treat the foot segment. For instance, we take the ankle and knee joint moments as the distal and proximal joint moments, respectively, for the shank segment. The knee and hip joint moments are then treated as the distal and proximal joint moments, respectively, for the thigh segment. Then, the knee and hip moments are calculated respectively as
${\boldsymbol{M}}_{P}^{S}=$ $-\boldsymbol{M}_{G}-\left(\boldsymbol{r}+l_{f}\boldsymbol{e}_{f}+l_{s}\boldsymbol{e}_{s}\right)\times\boldsymbol{F}_{G}+I_{f}\boldsymbol{\dot{\omega}}_{f}+I_{s}\boldsymbol{\dot{\omega}}_{s}-(l_{p}^{f}\boldsymbol{e}_{f}+$
$l_{s}\boldsymbol{e}_{s})\times m_{f}\left(\boldsymbol{a}_{f}+g\boldsymbol{e}_{z}\right)-l_{p}^{s}\boldsymbol{e}_{s}\times m_{s}\left(\boldsymbol{a}_{s}+g\boldsymbol{e}_{z}\right),$ (4)
@@
${\boldsymbol{M}}_{P}^{T}=$ $-({\boldsymbol{r}}+l_{f}{\boldsymbol{e}}_{f}+l_{s}{\boldsymbol{e}}_{s}+l_{t}{\boldsymbol{e}}_{t})\times{\boldsymbol{F}}_{G}-(l_{p}^{s}{\boldsymbol{e}}_{s}+l_{t}{\boldsymbol{e}}_{t})\times$
$m_{s}({\boldsymbol{a}}_{s}+g{\boldsymbol{e}}_{z})-\left(l_{p}^{f}\boldsymbol{e}_{f}+l_{s}\boldsymbol{e}_{s}+l_{t}\boldsymbol{e}_{t}\right)\times m_{f}\left(\boldsymbol{a}_{f}+g\boldsymbol{e}_{z}\right)$
$+I_{f}\boldsymbol{\dot{\omega}}_{f}+I_{s}\boldsymbol{\dot{\omega}}_{s}+I_{t}\boldsymbol{\dot{\omega}}_{t}-\boldsymbol{M}_{G},$ (5)
@@
The formulations (3), (4), and (5) are consistent for both the stance leg and swing leg. For the swing leg, the ground reaction forces and moments are zero, i.e., $\boldsymbol{F}_{G}=\boldsymbol{0}$ and $\boldsymbol{M}_{G}=\boldsymbol{0}$. Translation accelerations and angular accelerations of the leg segments are extracted from the optical marker measurements. The corresponding anthropometry information of the participant such as segment length, mass, and location of COM can be found in Chapter 4 of [25]. We used estimate ratios for the joint moment calculation in this study.